OpenStudy (anonymous):

Solve the system of the quadratic equation y=x^2-2x+2

OpenStudy (anonymous):

(x-2)(x-1)

OpenStudy (anonymous):

equal it to zero

OpenStudy (anonymous):

i would complete the square you will see two complex solutions

OpenStudy (anonymous):

\[x^2-2x+2=0\]subtract 2 from both sides \[x^2-2x=-2\] then take half of \(-2\) to get \(-1\) square it and get \(1\) and complete the square as \[(x-1)^2=-2+1=-1\]

OpenStudy (anonymous):

take the square root to get \[x-1=\pm i\] and so \[x=1\pm i\]